OpenStudy (anonymous):

whats the equation of the line that passes through the point (-2,1) and has slope 3 in point-slope form

OpenStudy (luigi0210):

Point-slope form: \[\LARGE y-y_{1}=m(x-x_{1})\] \[\LARGE (x_{1}, y_{1}) m=slope

OpenStudy (luigi0210):

*\[\LARGE (x_{1}, y_{1})\]

OpenStudy (anonymous):

(y-y1)/(x-x1)=m; (x,y) is any general point on the line ((y-1)/(x-2) =3 y-1 = 3x-6 y=3x-5

OpenStudy (anonymous):

welcome 2 os

OpenStudy (luigi0210):

Dang it lost ლ(ಠ_ಠლ)

OpenStudy (anonymous):

thanks. :)

OpenStudy (anonymous):

wut @Luigi0210